# Resumming Post-Minkowskian and Post-Newtonian gravitational waveform expansions

### v1

@@ 3.3 The upgoing solution
$C_{\rm up}={(2M)^{3}e^{-\partial_{\kappa_{1}}{\cal F}_{\rm inst}}\over x^{3+\frac{x}{2}}B_{++}}\left(1-{\textstyle\frac{B_{-+}B_{+-}}{B_{++}B_{--}}}\right)^{-1}=\frac{(2M)^{3}}{x^{3+\frac{x}{2}}}\frac{\Gamma(2-\ell)}{\Gamma(-2\ell)}+\ldots$ (65)
@@
4 Tests against MST and numerical simulations
In this Section we test our hypergeometric PM representation against the MST method Tagoshi:1993dm ; Poisson:1994yf ; Shibata:1994jx ; Tagoshi:1994sm ; Tanaka:1996lfd ; Mano:1996vt ; Mino:1997bx ; Suzuki:1998vy ; Fujita:2010xj ; Fujita:2011zk and numerical simulations. When we restrict ourselves to circular orbits [^3] As it is well known in this case the PM and PN expansion are just one expansion., we work at order 30PN, i.e. $v^{60}$, with $v$ the tangential velocity of the probe particle. We use the package Teukolsky of the Black Hole Perturbation Toolkit (BHPT) BHPToolkit both for the computations using the MST method and for numerical integration of the differential equation. The observables we consider are the $(\ell,m)$ harmonic modes of incoming solution $\mathfrak{R}_{\text{in}}$ and the energy flux radiated to infinity by the system. The results for the energy flux are also compared against the 22PN results in Fujita:2012cm .
@@ 4 Tests against MST and numerical simulations
4.1 The incoming and upgoing solutions
In this subsection we collect all the ingredients needed in the computations of the incoming, upgoing waveforms at order k=30 in the hypergeometric PM representation. We write444 Here we rewrite the second terms in the rhs of (60), (64), as the images under $a\to-a$ of the first terms, using the transformation rules $R_{-}\to R_{+}$, $\gamma\to\gamma^{-1}$, $F_{-\alpha^{\prime}}\leftrightarrow F_{+\alpha^{\prime}}$ and $B_{-\alpha^{\prime}}\leftrightarrow B_{+\alpha^{\prime}}$.
@@ 4.1 The incoming and upgoing solutions
$C_{\rm up}$ $=$ ${(2M)^{3}\over x^{3+\frac{x}{2}}\,\gamma_{\kappa}}{\Gamma({\textstyle\frac{5}{2}}-a+{\textstyle\frac{x}{2}})\over\Gamma(1-2a)}\left(1-e^{-2\pi{\rm i}a}\,{\textstyle\frac{\cos{\textstyle\frac{\pi}{2}}(x+2a)}{\cos{\textstyle\frac{\pi}{2}}(x-2a)}}\right)^{-1}\qquad,\qquad\gamma_{\kappa}=1+\sum_{i=1}^{\infty}\widehat{c}_{i,i-1}x^{i}$
@@
and [^5] We notice that $\gamma_{\rm tot}=e^{-\partial_{a}({\cal F}_{\rm tree}+{\cal F}_{\rm one-loop}+{\cal F}_{\rm inst})}$ collects the tree level, one-loop and instanton contributions to the prepotential of the underlying $SU(2)$ gauge theory with three fundamental flavours..
$\gamma_{\rm tot}=\underbrace{e^{-2\pi{\rm i}a}\,\frac{\Gamma({-}2a)^{2}\Gamma\left({-}\frac{3}{2}{+}a{-}\frac{x}{2}\right)\Gamma\left(\frac{1}{2}{+}a{-}\frac{x}{2}\right)\Gamma\left(\frac{5}{2}{+}a{-}\frac{x}{2}\right)}{\Gamma(2a)^{2}\Gamma\left(-\frac{3}{2}{-}a{-}\frac{x}{2}\right)\Gamma\left(\frac{1}{2}{-}a{-}\frac{x}{2}\right)\Gamma\left(\frac{5}{2}{-}a{-}\frac{x}{2}\right)}}_{\gamma_{0}=\frac{F_{-+}B_{+-}}{F_{++}B_{--}}}\underbrace{x^{2a}e^{-\partial_{a}{\cal F}_{\rm inst}}}_{\gamma}~{},~{}{\cal F}_{\rm inst}=-\sum_{i=1}^{\infty}{u_{i}\over i}x^{i}$ (69)
@@
$P_{0}(y)=1+\sum_{i=1}^{k}\sum_{j=0}^{i-1}c_{ij}\,x^{i}y^{j-i}\qquad,\qquad\widehat{P}_{0}(y)=\sum_{i=1}^{k}\sum_{j=0}^{i-1}\widehat{c}_{ij}\,x^{i}y^{j-i}\$ (70)
@@
$A_{1}$ $=$ $-8\left(\Delta_{ij}+1\right)\left(\Delta_{ij}-x-5\right)~{},~{}\widehat{A}_{1}=4\left(4a^{2}-(x+5)^{2}\right)\Delta_{i,j}$
$A_{3}$ $=$ $8\Delta_{ij}\left(\Delta_{ij}-x-5\right)~{},~{}\widehat{A}_{3}=4\left(4a^{2}-(x+5)^{2}\right)\left(-\Delta_{i,j}+x-2\right)$
@@
$B_{1}$ $=$ $-16\left(\Delta_{ij}+1\right)~{},~{}\widehat{B}_{1}=8\Delta_{ij}\left(\Delta_{ij}+x+5\right)~{},~{}\widehat{B}_{3}=8\left(-\Delta_{i,j}^{2}-7\Delta_{ij}+x^{2}+3x-10\right)$
$B_{2}$ $=$ $4\left(-2(x-1)\Delta_{i,j}+4a^{2}+x^{2}-2x-3\right)~{},~{}B_{3}=16\Delta_{ij}$
@@
$C_{1}$ $=$ $8\left(\Delta_{ij}-x-5\right)~{},~{}\widehat{C}_{1}=16a^{2}-4(x+5)^{2}~{},~{}C_{2}=16~{},~{}\widehat{C}_{2}=8\left(\Delta_{ij}+x+5\right)$ (72)
@@
Inverting the series at high PM order can be time consuming for a Mathematica code. There is an alternative, more direct way to find the function $a(u)$ based on the connection between this quantity and the quantum SW period of a $SU(2)$ theory with three fundamental flavours. The quantum SW period $a$ is known to satisfy the continuous fraction equation Poghosyan:2020zzg
@@
$P(a)=a^{2}+ax+2x-\frac{3x^{2}}{4}x-(\ell+{\textstyle\frac{1}{2}})^{2}\qquad,\qquad M(a)=(a-{\textstyle\frac{x}{2}}-{\textstyle\frac{1}{2}})(a-{\textstyle\frac{x}{2}}+{\textstyle\frac{3}{2}})(a+{\textstyle\frac{x}{2}}+{\textstyle\frac{3}{2}})$ (75)
@@ 4.2 The luminosity
$b^{0}_{\ell m}$ $=$ ${\pi\over 2}\sqrt{(\ell-1)\ell(\ell+1)(\ell+2)}\,Y_{0}^{\ell m}({\textstyle\frac{\pi}{2}},0)$
$b^{1}_{\ell m}$ $=$ $\pi\sqrt{(\ell-1)(\ell+2)}Y_{-1}^{\ell m}({\textstyle\frac{\pi}{2}},0)$
$b^{2}_{\ell m}$ $=$ $\pi\,Y_{-2}^{\ell m}({\textstyle\frac{\pi}{2}},0),$ (79)
@@
the energy and angular momentum of the light particle. Finally, the luminosity, i.e. the amount of energy per unit time emitted by the system towards infinity, is computed by the formula (see Mino:1997bx and references therein)
@@
$\left({dE\over dt}\right)_{N}=\frac{32\mu^{2}v^{10}}{5M^{2}}=\mu^{2}\,\eta_{22}^{0}$ (83)
@@ 4.3 Numerical tests
The results are compared in Table 1 against those obtained by the MST method and numerical integration using the BHPT BHPToolkit . In all the subsequent tables, the displayed digits are those ones that are insensible to the contribution of terms of order higher than $v^{60}$ in $P_{0}(y)$ and $\widehat{P}_{0}(y)$. The agreement is up to 10 and 14 digits even for small values of $r$, where the PM approximation is less reliable.
@@
Finally in Tables 3 and 4 we compare our 30PN results for the energy loss against those obtained by MST/Numerical and by the 22PN computation in Fujita:2012cm .
@@ 5 The Heun equation
${\rm Heun}_{-}(z)=1+\ldots\qquad,\qquad{\rm Heun}_{+}(z)=z^{1-\gamma_{0}}(1+\ldots)$ (86)
@@
$\beta_{0}=\kappa_{1}{+}\kappa_{3}\quad,\quad\beta_{1}=\kappa_{1}{+}\kappa_{2}\quad,\quad\gamma_{1}=\kappa_{2}{+}\kappa_{3}\quad,\quad\sigma=\kappa_{1}^{2}-\widehat{u}\quad,\quad\gamma_{2}=1{-}\gamma_{0}+2\kappa_{1}$ (87)
@@
$G(z)$ $\underset{z\to 0}{\approx}$ $B_{-}(1+\ldots)+B_{+}z^{1-\gamma_{0}}(1+\ldots)$
@@ 5.1 The hypergeometric representation
$P_{0}(y)$ $=$ $1+\sum_{i=1}^{k}\sum_{j=0}^{i-1}c_{ij}\,x^{i}y^{j-i}\qquad,\qquad\widehat{P}_{0}(y)=\sum_{i=1}^{k}\sum_{j=0}^{i}\widehat{c}_{ij}\,x^{i}y^{j-i}$
$P_{1}(z)$ $=$ $1+\sum_{j=1}^{k}\sum_{i=0}^{j-1}d_{ij}\,z^{i-j}x^{j}\qquad,\qquad\widehat{P}_{1}(z)=\sum_{j=1}^{k}\sum_{i=0}^{j}\widehat{d}_{ij}\,z^{i-j}x^{j}$ (92)
@@
$\widehat{c}_{11}$ $=$ $-\widehat{c}_{10}=\frac{\left(2a^{2}-\kappa_{2}-\kappa_{3}+2\kappa_{2}\kappa_{3}\right)}{4a^{2}{-}1}$
$d_{01}$ $=$ $1-\frac{y\left(2\gamma_{0}{-}2\kappa_{1}{-}1\right)\left(a^{2}{-}\kappa_{1}^{2}\right)}{4a^{2}-1}$
$\widehat{d}_{11}$ $=$ $-\widehat{d}_{01}=\frac{\left(2\kappa_{1}^{2}{+}2a^{2}{-}2\gamma_{0}\kappa_{1}{+}\gamma_{0}{-}1\right)}{4a^{2}{-}1}$ (93)
@@
$\widehat{u}_{1}$ $=$ $\frac{\gamma_{0}\left[a^{2}\left(1-2\kappa_{1}-2\kappa_{2}-2\kappa_{3}\right)+\kappa_{1}\kappa_{2}+\kappa_{1}\kappa_{3}+\kappa_{2}\kappa_{3}-2\kappa_{1}\kappa_{2}\kappa_{3}\right]}{4a^{2}-1}$ (94)
$-\frac{\left(a^{2}-\kappa_{1}^{2}\right)\left(2a^{2}-\kappa_{2}+2\kappa_{2}\kappa_{3}-\kappa_{3}\right)}{4a^{2}-1}$
@@ 5.2 Connection formulae
$H^{0}_{\alpha}(y)$ $=$ $\sum_{\alpha^{\prime}=\pm}B_{\alpha\alpha^{\prime}}\widetilde{H}^{0}_{\alpha^{\prime}}(y)\qquad,\qquad H^{1}_{\alpha}(z)=\sum_{\alpha^{\prime}=\pm}F_{\alpha\alpha^{\prime}}\widetilde{H}^{1}_{\alpha^{\prime}}(z)$ (98)
@@
Plugging (98) into (5.1) one finds the connection formulae
$G^{0}_{\alpha}(y)$ $=$ $\sum_{\alpha^{\prime}=\pm}B_{\alpha\alpha^{\prime}}\widetilde{G}^{0}_{\alpha^{\prime}}(y)\qquad,\qquad G^{1}_{\alpha}(z)=\sum_{\alpha^{\prime}=\pm}F_{\alpha\alpha^{\prime}}\widetilde{G}^{1}_{\alpha^{\prime}}(z)$ (100)

### v2

@@ III.3 The upgoing solution
$C_{\rm up}$ $=$ ${(2M)^{3}e^{-\partial_{\kappa_{1}}{\cal F}_{\rm inst}}\over x^{3+\frac{x}{2}}B_{++}}\left(1-{\textstyle\frac{B_{-+}B_{+-}}{B_{++}B_{--}}}\right)^{-1}=$ (67)
$=\frac{(2M)^{3}}{x^{3+\frac{x}{2}}}\frac{\Gamma(2-\ell)}{\Gamma(-2\ell)}+\ldots$
@@
IV Tests against MST and numerical simulations
In this Section we test our hypergeometric PM representation against the MST method [10, 11, 12, 13, 14, 15, 16, 17, 18, 19] and numerical simulations. When we restrict ourselves to circular orbits [^2] As it is well known in this case the PM and PN expansion are just one expansion., we work at order 30PN, i.e. $v^{60}$, with $v$ the tangential velocity of the probe particle. We use the package Teukolsky of the Black Hole Perturbation Toolkit (BHPT) [56] both for the computations using the MST method and for numerical integration of the differential equation. The observables we consider are the $(\ell,m)$ harmonic modes of incoming solution $\mathfrak{R}_{\text{in}}$ and the energy flux radiated to infinity by the system. The results for the energy flux are also compared against the 22PN results in [20].
@@ IV Tests against MST and numerical simulations
IV.1 The incoming and upgoing solutions
In this subsection we collect all the ingredients needed in the computations of the incoming, upgoing waveforms at order k=30 in the hypergeometric PM representation. We write333 Here we rewrite the second terms in the right hand sides of (62), (66), as the images under $a\to-a$ of the first terms, using the transformation rules $R_{-}\to R_{+}$, $\gamma\to\gamma^{-1}$, $F_{-\alpha^{\prime}}\leftrightarrow F_{+\alpha^{\prime}}$ and $B_{-\alpha^{\prime}}\leftrightarrow B_{+\alpha^{\prime}}$.
@@ IV.1 The incoming and upgoing solutions
$C_{\rm up}$ $=$ ${(2M)^{3}\over x^{3+\frac{x}{2}}\,\gamma_{\kappa}}{\Gamma({\textstyle\frac{5}{2}}-a+{\textstyle\frac{x}{2}})\over\Gamma(1-2a)}\left(1-e^{-2\pi{\rm i}a}\,{\textstyle\frac{\cos{\textstyle\frac{\pi}{2}}(x+2a)}{\cos{\textstyle\frac{\pi}{2}}(x-2a)}}\right)^{-1}$
$\gamma_{\kappa}$ $=$ $1+\sum_{i=1}^{\infty}\widehat{c}_{i,i-1}x^{i}$
@@
and [^4] We notice that $\gamma_{\rm tot}=e^{-\partial_{a}({\cal F}_{\rm tree}+{\cal F}_{\rm one-loop}+{\cal F}_{\rm inst})}$ collects the tree level, one-loop and instanton contributions to the prepotential of the underlying $SU(2)$ gauge theory with three fundamental flavours.
$\gamma_{\rm tot}=\underbrace{e^{-2\pi{\rm i}a}\,\frac{\Gamma({-}2a)^{2}\Gamma\left({-}\frac{3}{2}{+}a{-}\frac{x}{2}\right)\Gamma\left(\frac{1}{2}{+}a{-}\frac{x}{2}\right)\Gamma\left(\frac{5}{2}{+}a{-}\frac{x}{2}\right)}{\Gamma(2a)^{2}\Gamma\left(-\frac{3}{2}{-}a{-}\frac{x}{2}\right)\Gamma\left(\frac{1}{2}{-}a{-}\frac{x}{2}\right)\Gamma\left(\frac{5}{2}{-}a{-}\frac{x}{2}\right)}}_{\gamma_{0}=\frac{F_{-+}B_{+-}}{F_{++}B_{--}}}\underbrace{x^{2a}e^{-\partial_{a}{\cal F}_{\rm inst}}}_{\gamma}\leavevmode\nobreak\ ,\leavevmode\nobreak\ {\cal F}_{\rm inst}=-\sum_{i=1}^{\infty}{u_{i}\over i}x^{i}$ (71)
@@
$P_{0}(y)=1+\sum_{i=1}^{k}\sum_{j=0}^{i-1}c_{ij}\,x^{i}y^{j-i}\,,\,\widehat{P}_{0}(y)=\sum_{i=1}^{k}\sum_{j=0}^{i-1}\widehat{c}_{ij}\,x^{i}y^{j-i}\$ (72)
@@
$A_{1}$ $=$ $-8\left(\Delta_{ij}+1\right)\left(\Delta_{ij}-x-5\right)\leavevmode\nobreak\ ,\leavevmode\nobreak\ \widehat{A}_{1}=4\left(4a^{2}-(x+5)^{2}\right)\Delta_{i,j}$
$A_{3}$ $=$ $8\Delta_{ij}\left(\Delta_{ij}-x-5\right)\leavevmode\nobreak\ ,\leavevmode\nobreak\ \widehat{A}_{3}=4\left(4a^{2}-(x+5)^{2}\right)\left(-\Delta_{i,j}+x-2\right)$
@@
$B_{1}$ $=$ $-16\left(\Delta_{ij}+1\right)\leavevmode\nobreak\ ,\leavevmode\nobreak\ \widehat{B}_{1}=8\Delta_{ij}\left(\Delta_{ij}+x+5\right)\leavevmode\nobreak\ ,\leavevmode\nobreak\ \widehat{B}_{3}=8\left(-\Delta_{i,j}^{2}-7\Delta_{ij}+x^{2}+3x-10\right)$
$B_{2}$ $=$ $4\left(-2(x-1)\Delta_{i,j}+4a^{2}+x^{2}-2x-3\right)\leavevmode\nobreak\ ,\leavevmode\nobreak\ B_{3}=16\Delta_{ij}$
@@
$C_{1}$ $=$ $8\left(\Delta_{ij}-x-5\right)\leavevmode\nobreak\ ,\leavevmode\nobreak\ \widehat{C}_{1}=16a^{2}-4(x+5)^{2}\leavevmode\nobreak\ ,\leavevmode\nobreak\ C_{2}=16\leavevmode\nobreak\ ,\leavevmode\nobreak\ \widehat{C}_{2}=8\left(\Delta_{ij}+x+5\right)$ (74)
@@
Inverting the series at high PM order can be time consuming for a Mathematica code. There is an alternative, more direct way to find the function $a(u)$ based on the connection between this quantity and the quantum SW period of a $SU(2)$ theory with three fundamental flavours. The quantum SW period $a$ is known to satisfy the continuous fraction equation [57]
@@
$P(a)$ $=$ $a^{2}+ax+2x-\frac{3x^{2}}{4}x-(\ell+{\textstyle\frac{1}{2}})^{2},$
$M(a)$ $=$ $(a-{\textstyle\frac{x}{2}}-{\textstyle\frac{1}{2}})(a-{\textstyle\frac{x}{2}}+{\textstyle\frac{3}{2}})(a+{\textstyle\frac{x}{2}}+{\textstyle\frac{3}{2}})$ (77)
@@ IV.2 The luminosity
$b^{0}_{\ell m}={\pi\over 2}\sqrt{(\ell-1)\ell(\ell+1)(\ell+2)}\,Y_{0}^{\ell m}({\textstyle\frac{\pi}{2}},0),\,b^{1}_{\ell m}=\pi\sqrt{(\ell-1)(\ell+2)}Y_{-1}^{\ell m}({\textstyle\frac{\pi}{2}},0),\,b^{2}_{\ell m}=\pi\,Y_{-2}^{\ell m}({\textstyle\frac{\pi}{2}},0),$ (81)
and
@@
the energy and angular momentum of the light particle. Finally, the luminosity, i.e. the amount of energy per unit time emitted by the system towards infinity, is computed by the formula (see [16] and references therein)
@@
$\left({d{\cal E}\over dt}\right)_{N}=\frac{32\mu^{2}v^{10}}{5M^{2}}=\mu^{2}\,\eta_{22}^{0}$ (85)
@@ IV.3 Numerical tests
The results are compared in Table 1 against those obtained by the MST method and numerical integration using the BHPT [56]. In all the subsequent tables, the displayed digits are those ones that are insensible to the contribution of terms of order higher than $v^{60}$ in $P_{0}(y)$ and $\widehat{P}_{0}(y)$. The agreement is up to 10 and 14 digits even for small values of $r$, where the PM approximation is less reliable.
@@
Finally in Tables 3 and 4 we compare our 30PN results for the energy loss against those obtained by MST/Numerical and by the 22PN computation in [20].
@@ V The Heun equation
${\rm Heun}_{-}(z)=1+\ldots\,,\,{\rm Heun}_{+}(z)=z^{1-\gamma_{0}}(1+\ldots)$ (88)
@@
$\beta_{0}$ $=$ $\kappa_{1}{+}\kappa_{3}\,,\,\beta_{1}=\kappa_{1}{+}\kappa_{2}\,,\,\sigma=\kappa_{1}^{2}-\widehat{u},$
$\gamma_{1}$ $=$ $\kappa_{2}{+}\kappa_{3}\,,\,\gamma_{2}=1{-}\gamma_{0}+2\kappa_{1}$ (89)
@@
$G(z)$ $\underset{z\to 1}{\approx}$ $D_{-}(1+\ldots)+D_{+}(1{-}z)^{1-\kappa_{2}-\kappa_{3}}(1+\ldots)$
$G(z)$ $\underset{z\to 0}{\approx}$ $B_{-}(1+\ldots)+B_{+}z^{1-\gamma_{0}}(1+\ldots)$ (91)
@@ V.1 The hypergeometric representation
$P_{0}(y)$ $=$ $1+\sum_{i=1}^{k}\sum_{j=0}^{i-1}c_{ij}\,x^{i}y^{j-i}\,,\,\widehat{P}_{0}(y)=\sum_{i=1}^{k}\sum_{j=0}^{i}\widehat{c}_{ij}\,x^{i}y^{j-i}$
$P_{1}(z)$ $=$ $1+\sum_{j=1}^{k}\sum_{i=0}^{j-1}d_{ij}\,z^{i-j}x^{j}\,,\,\widehat{P}_{1}(z)=\sum_{j=1}^{k}\sum_{i=0}^{j}\widehat{d}_{ij}\,z^{i-j}x^{j}$
@@
$\widehat{c}_{11}=$ $-\widehat{c}_{10}=\frac{\left(2a^{2}-\kappa_{2}-\kappa_{3}+2\kappa_{2}\kappa_{3}\right)}{4a^{2}{-}1}\qquad\qquad\qquad$
$d_{01}=$ $1-\frac{y\left(2\gamma_{0}{-}2\kappa_{1}{-}1\right)\left(a^{2}{-}\kappa_{1}^{2}\right)}{4a^{2}-1}\qquad\qquad\qquad\quad$
$\widehat{d}_{11}=$ $-\widehat{d}_{01}=\frac{\left(2\kappa_{1}^{2}{+}2a^{2}{-}2\gamma_{0}\kappa_{1}{+}\gamma_{0}{-}1\right)}{4a^{2}{-}1}\qquad\qquad\quad$
@@
$\widehat{u}_{1}=\frac{\gamma_{0}\left[a^{2}\left(1-2\kappa_{1}-2\kappa_{2}-2\kappa_{3}\right)+\kappa_{1}\kappa_{2}+\kappa_{1}\kappa_{3}+\kappa_{2}\kappa_{3}-2\kappa_{1}\kappa_{2}\kappa_{3}\right]}{4a^{2}-1}-\frac{\left(a^{2}-\kappa_{1}^{2}\right)\left(2a^{2}-\kappa_{2}+2\kappa_{2}\kappa_{3}-\kappa_{3}\right)}{4a^{2}-1}\$ (95)
@@ V.2 Connection formulae
$H^{0}_{\alpha}(y)=\sum_{\alpha^{\prime}=\pm}B_{\alpha\alpha^{\prime}}\widetilde{H}^{0}_{\alpha^{\prime}}(y),H^{1}_{\alpha}(z)=\sum_{\alpha^{\prime}=\pm}F_{\alpha\alpha^{\prime}}\widetilde{H}^{1}_{\alpha^{\prime}}(z)$ (99)
@@
Plugging (99) into (V.1) one finds the connection formulae
$G^{0}_{\alpha}(y)$ $=$ $\sum_{\alpha^{\prime}=\pm}B_{\alpha\alpha^{\prime}}\widetilde{G}^{0}_{\alpha^{\prime}}(y),$